OpenStudy (anonymous):

The expression below is a Riemann Sum approximation for which of the following? (1/50)(sqrt(1/50)+sqrt(2/50)+sqrt(3/50)...sqrt(50/50)) A. integral 0S1 sqrt(x/50) dx B. integral 0S1 sqrt(x) dx C. (1/50) integral 0S1 sqrt(x)dx

OpenStudy (kinggeorge):

I saw this same question on here a few days ago... Anyways, since you're going to \(\sqrt{50\over50}=\sqrt{1}\) and each interval is exactly \(1/50\), it's the integral from 0 to 1 of \(\sqrt{x}\)